Circular No. 6148

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CIRCINI 1995 A. Evans, Keele University; and R. V. Yudin, Pulkovo Observatory, re-

port observations of this nova with the 0.75-m telescope (+ UCT polarimeter) at the South African Astronomical Observatory during Jan. 31–Feb. 5: “We find a polarization through Johnson V of 3.7 ± 0.3 percent in p.a. 55^{o} ± 8^{o} (typical errors cited), with no evidence for variability. These values are broadly similar to those of field stars, suggesting that the polarization of the nova is interstellar. These observations imply that the reddening may be high [E(B − V) > 0.5].”

CH CYGNI B. R. Espey and G. A. Kriss, Hopkins Ultraviolet Telescope (HUT)

team, Astro-2 Space Shuttle mission; and J. J. Johnson and R. E. Schulte-Ladbeck, Wisconsin Ultraviolet Photopolarimeter Experiment (WUPPE) team, report ultraviolet and far-ultraviolet observations of CH Cyg on Mar. 8.22 UT with both the HUT (range 80.0–180.0 nm, resolution ∼ 0.3 nm) and WUPPE (range 140.0–320.0 nm, resolution ∼ 1.6 nm), 7 days after the ejection event noted by Iijima (IAUC 6146). Preliminary results follow: “The ultraviolet spectrum is generally similar to that seen during previous ejection events. The strongest emission lines seen are those of C II (133.4 nm), O I] (135.6 nm), Si IV + O IV] (140.0 nm), N IV] (148.6 nm), C IV (154.9 nm), Ni II (159.6 nm), O III] (166.3 nm), N I (174.2 nm), N III] (174.8 nm), Si III] (189.3 nm), C III] (190.9 nm), C II] (232.6 nm), and Mg II (279.8 nm). Lyα and O I 130.4-nm are hidden in the geocoronal emission profiles. In addition to these features seen in previous IUE spectra, HUT has detected emission from C III at 97.7 nm, O VI at 103.2 and 103.8 nm, N II + He II at 108.5 nm, and C III at 117.6 nm. P-Cyg profiles are in evidence for the C III 117.6-nm, Si IV 140.0-nm, He II 164.0-nm, and N I 174.2-nm lines, with weaker absorption to the blue of the C IV 154.9-nm line. The typical width of the ultraviolet absorption profiles is ∼ 2000 km/s, slightly broader than those seen in the optical observations reported by Iijima (ibid.). Weaker lines in evidence include the Fe II ‘UV191 transitions’ at 178.6 and 250.9 nm, which are pumped by Lyα radiation. No emission from Si II was detected. The continuum rises to the red, with evidence for Fe II absorption, from 2 × 10^{−13} erg cm^{−2} s^{−1} A^{−1} at 100.0 nm to 1 × 10^{−12} erg cm^{−2} s^{−1} A^{−1} at 180.0 nm. After a dip to about half this value around 235.0 nm, the flux recovers towards 320.0 nm.”

1995 March 11 Daniel W. E. Green